OpenStudy (anonymous):

(x+4)+2/(x+1) < 0,solve for x . very very urgent @terenzreignz

OpenStudy (anonymous):

pls help me

terenzreignz (terenzreignz):

Is it this: \[\Large (x+4)+\frac{2}{x+1}< 0 \]?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, first, you need to combine them into one fraction.

OpenStudy (anonymous):

listen its coming (x+3) (x+2) (x+1)<0,then how to find the range of x

terenzreignz (terenzreignz):

What's coming? Anyway, that's rather simple, you want the product (x+3)(x+2)(x+1) to be negative so you need either one or three of these factors to be negative.

OpenStudy (anonymous):

listen the answer given is -2<x<-1 and -3<x,how?????????????

terenzreignz (terenzreignz):

lol okay, the first step is to find the critical numbers of (x+3)(x+2)(x+1) Critical numbers are values of x which would make that product zero.

OpenStudy (anonymous):

x=-1,-2 -3

terenzreignz (terenzreignz):

That's right :) Now, let's make a table \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&&&&&&&\\x+2&&&&&&&\\x+1&&&&&&&\\(x+3)(x+2)(x+1)\end{matrix}\right.\]

terenzreignz (terenzreignz):

OKay? Now, what is the critical number associated with x+3?

OpenStudy (anonymous):

infinity

terenzreignz (terenzreignz):

No, I mean, when is x + 3 equal to zero, when x = -3, right?

OpenStudy (anonymous):

yap

terenzreignz (terenzreignz):

Okay, so it's zero when x = -3 \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&&\color{red}0&&&&&\\x+2&&&&&&&\\x+1&&&&&&&\\(x+3)(x+2)(x+1)\end{matrix}\right.\]

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, before -3, what values does x+3 take? Positive or negative

OpenStudy (anonymous):

negative

terenzreignz (terenzreignz):

that's right :) \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&&&&&&&\\x+1&&&&&&&\\(x+3)(x+2)(x+1)\end{matrix}\right.\] And positive after -3. Next, what's the critical number for x+2?

OpenStudy (anonymous):

0

terenzreignz (terenzreignz):

No, I mean, when is x+2 = 0 ?

OpenStudy (anonymous):

x=-2

terenzreignz (terenzreignz):

Right. So it's zero here \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&&&&\color{red}0&&&\\x+1&&&&&&&\\(x+3)(x+2)(x+1)\end{matrix}\right.\] To the left, is it positive or negative?

OpenStudy (anonymous):

positive but i am still confused

terenzreignz (terenzreignz):

No... I mean, for values less than -2, then x + 2 would be negative, right? Try -3 -3 + 2 = -1 which is negative. Try -2.5 -2.5 + 2 = -0.5 which is negative. It's negative for all values of x less than -2

terenzreignz (terenzreignz):

\[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&-&-&-&\color{red}0&+&+&+\\x+1&&&&&&&\\(x+3)(x+2)(x+1)\end{matrix}\right.\]

OpenStudy (anonymous):

-2<x<-1 and -3<x,how?????????????

terenzreignz (terenzreignz):

Yes, we're getting there... :) Now just do the last row, when is x+1 equal to zero?

OpenStudy (anonymous):

x=-1

terenzreignz (terenzreignz):

Right, and it's negative for all values of x less than -1, etc \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&-&-&-&\color{red}0&+&+&+\\x+1&-&-&-&-&-&\color{red}0&+\\(x+3)(x+2)(x+1)\end{matrix}\right.\]

OpenStudy (anonymous):

so

terenzreignz (terenzreignz):

Well, over here, we simply have the product of the three, in this row: \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&-&-&-&\color{red}0&+&+&+\\x+1&-&-&-&-&-&\color{red}0&+\\(x+3)(x+2)(x+1)&?&?&?&?&?&?&?\end{matrix}\right.\]

OpenStudy (anonymous):

then

terenzreignz (terenzreignz):

We just multiply signs LOL Over here in the first column \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&-&-&-&\color{red}0&+&+&+\\x+1&-&-&-&-&-&\color{red}0&+\\(x+3)(x+2)(x+1)&\color{blue}?&?&?&?&?&?&?\end{matrix}\right.\] We have three negative signs... what's negative times negative times negative?

OpenStudy (anonymous):

thanx

terenzreignz (terenzreignz):

We're actually not yet done, but if you think you can take it from here, then, be my guest :P

OpenStudy (anonymous):

no,explain it to me

terenzreignz (terenzreignz):

First, what is negative x negative x negative? Is it positive or negative?

OpenStudy (anonymous):

-

terenzreignz (terenzreignz):

That's right. \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&-&-&-&\color{red}0&+&+&+\\x+1&-&-&-&-&-&\color{red}0&+\\(x+3)(x+2)(x+1)&-&?&?&?&?&?&?\end{matrix}\right.\]

terenzreignz (terenzreignz):

Now, here, it's zero, because one of the factors is zero: \[\Large \left.\begin{matrix}&-\infty&-3&&-2&&-1&\infty\\x+3&-&\color{red}0&+&+&+&+&+\\x+2&-&-&-&\color{red}0&+&+&+\\x+1&-&-&-&-&-&\color{red}0&+\\(x+3)(x+2)(x+1)&-&\color{red}0&\color{blue}?&?&?&?&?\end{matrix}\right.\] What about there^? (third column)

OpenStudy (anonymous):

ok i got it.bye

terenzreignz (terenzreignz):

Oh... okay then :)